OpenStudy (anonymous):

The Washington Monument is 170 meters tall. How long would it take for a penny to fall?

OpenStudy (anonymous):

So far i have distance which equals 170m, and i am guessing that this person is dropping the coin so it would mean that it is falling from rest which gives us initial velocity 0, and we have acceleration which equals gravity which equals -9.8m/s^2. I am going to use the equation which has no Vf (i don't know what it exactly means, so if you know it would be great if you could tell me).

OpenStudy (anonymous):

The equation i am going to use is d = Vo(t) + 1/2 (-9.8)(t)^2.

OpenStudy (anonymous):

So far i have put down -170m = 1/2(-9.8)(t)^2. This is where i kind of get lost. My teacher says that to get a positive answer we should make the distance negative, but she said it isn't mandatory. From there, she got the answer 5.89 or something around that, and i just don't know how she did that.

OpenStudy (anonymous):

Can you please show me step by step how it is done? Thank you

OpenStudy (anonymous):

@Nurali , @Luigi0210 , @blues can any of you guys please help me?

OpenStudy (anonymous):

@SolomonZelman @robtobey

OpenStudy (wach):

I can try!

OpenStudy (wach):

I think your preliminary equation is correct. The formula, x=xo+vot+1/2at(t^2) relates the distance of an object to its force. Let's map out we know: x = -170 m(because it's in the negative direction - we're taking the top of the monument to be equal to 0m, do you see how it's negative?) vo = velocity initial = 0 a= -9.8 m/s^s = gravity t = unknown variable

OpenStudy (wach):

From here, we just solve for the time it takes, right? -170 = 0.5(-9.8)(t^2)

OpenStudy (wach):

-170 = 4.9t^2 divide both sides 34.69 = t^2 now, take the square root of t^2 to get only t, we want to remove the ^2 part t=5.89 seconds this is how long it takes for the penny to reach the ground

OpenStudy (anonymous):

thank you, I realized my mistake and i understand how to solve the problem :D

OpenStudy (wach):

No problem :)